# Analysis of Failures and Risks in Deep Learning Model Converters: A Case Study in the ONNX Ecosystem

### v3

@@ Abstract.
This paper analyzes failures in DL model converters. We survey software engineers about DL interoperability tools, use cases, and pain points (N=92). Then, we characterize failures in model converters associated with the main interoperability tool, ONNX (N=200 issues in PyTorch and TensorFlow). Finally, we formulate and test two hypotheses about structural causes for the failures we studied. We find that the node conversion stage of a model converter accounts for 75% of the defects and 33% of reported failure are related to semantically incorrect models. The cause of semantically incorrect models is elusive, but models with behaviour inconsistencies share operator sequences. Our results motivate future research on making DL interoperability software simpler to maintain, extend, and validate. Research into behavioural tolerances and architectural coverage metrics would be fruitful.
@@ 1. Introduction
Deep Learning (DL) achieves state-of-the-art performance in many domains (Grigorescu et al. 2020; Kim et al. 2019). Software engineers engage in many activities for deep learning, including developing, re-using, fine-tuning, and deploying DL models (Amershi et al. 2019; Han et al. 2021; Jiang et al. 2023c, a). They use tools at each stage: DL frameworks for development (e.g., PyTorch (Paszke et al. 2019)); DL model registries for re-use (e.g., HuggingFace (Face 2021)), and DL compilers for deployment platforms (e.g., TVM (Chen et al. 2018)). Preferably, these tools would be interoperable, so that DL models can move seamlessly from one to another. Model conversion errors disrupt engineering workflows or compromise the resulting models (Openja et al. 2022). High-quality model converters are crucial to the deep learning ecosystem.
Researchers have characterized failures in most of the DL ecosystem, but not in model converters. As depicted in Figure 1, previous works have considered DL development frameworks (Chen et al. 2022; Long and Chen 2022; Islam et al. 2019a; Pham et al. 2020; Tan et al. 2022) and DL deployment compilers and runtimes (Shen et al. 2021a; Guo et al. 2019). In contrast, prior research on DL model converters is limited to measuring conversions of 5 DL models (Openja et al. 2022). We lack systematic knowledge of failure symptoms, causes, and patterns in DL model converters.
In this work, we analyze failures of DL model converters. We first survey software engineers (N=108), focusing on their experiences with DL interoperability tools, their use cases, and failure modes encountered. Then we analyze failures in DL model converters to the ONNX IR (Open Neural Network eXchange’s Intermediate Representation), the most prominent interoperability target for DL models. We sample 200 closed GitHub issues (100 per converter) and determine failure symptoms, causes, and locations. Finally, we examine two possible root causes of model converter failures, testing hypotheses about specification updates and model types.
@@ 2.1. DL Model Conversion as Interoperability
In the context of DL, interoperability focuses on model reuse and is defined as the ability of DL software to exchange DL models (i.e., deep neural networks/DNNs) (Liu et al. 2020). Wegner describes two patterns for interoperability: creating pairwise mappings between systems; and introducing a common intermediary understood by all participants (Wegner 1996). These patterns are illustrated in Figure 1 in the context of DL. The two patterns trade-off customizability for scalability.
Two kinds of DL systems interoperate on DL models: frameworks for DL model development, and runtimes for DL model deployment. In the downward path of Figure 1, pairwise mappings occur in DL compilers such as TVM (Chen et al. 2018), which map from framework representations into internal compiler representations for each supported framework. Along the rightward path, common intermediaries such as ONNX (ONN 2019) and MMDnn (Liu et al. 2020) give standard representations for DL models. In this pattern, each framework and runtime has a one- or two-way adapter to the common intermediary.
DL Model Converters: Model converters fill a purpose similar to compiler front-ends (Li et al. 2020). They transform a model from a DL framework into a high-level IR representing the model’s computations and control flow. Graph-level optimizations are applied before further conversion to a low-level IR for hardware optimization and code generation. Table 1 summarizes a typical design.
@@
Figure 2 illustrates a model converted from PyTorch to ONNX. Conversion is challenging, as noted by AirBus (ONNX 2023; Gauffriau and Pagetti 2023) and others (Consortium 2023), because it maps between graphs expressed with different operators and different semantics. Model conversion can produce models that are incompatible with runtimes or have different behaviours. For a compatibility issue, in PyTorch #78721 a converted ONNX model had a type mismatch (vbogach 2022). For a behavioral issue, in PyTorch #74732 a converted ONNX model’s prediction changed (YsYusaito 2022).
After conversion to an IR, models can be rendered back to DL frameworks or deployed to hardware. Framework-to-framework conversion can bypass issues in model reengineering (Jiang et al. 2023b). Deployment from an IR allows DL runtimes to optimize against the IR rather than the many DL frameworks.
@@ 2.2. Failure Studies of DL Components
Software engineering failures inform development and maintenance (Petroski et al. 1994; Anandayuvaraj and Davis 2022; Amusuo et al. 2022a). Previous failure analyses of DL interoperability software focused on the “Development” and “Runtime” components of Figure 1. Chen et al. studied 800 defects from 4 DL frameworks to obtain testing guidelines (Chen et al. 2022). Shen et al. studied DL compiler defects (Shen et al. 2021a), and others have tested DL compilers (Liu et al. 2023; Xiao et al. 2022). Some of this work has incidentally examined interoperability failures (Liu et al. 2023; Morovati et al. 2024); for example, Shen et al.’s study of DL compilers included “front ends” offering interoperability through pairwise (DL framework specific) and common intermediary (ONNX) model loaders (Shen et al. 2021a);
The two prior failure studies of DL model converters focused on testing and fault localization. Openja et al. converted 5 popular DL models (Openja et al. 2022). They considered only the current state of model converters, not past failures. Louloudakis et al. studied behavioral issues resulting from framework-to-framework conversion (Louloudakis et al. 2023a). They found failures in 10 out of 36 conversions. They created a fault localization and repair pipeline to localize and fix discrepancies (Louloudakis et al. 2023b).
@@ 3.2. Study Design
Figure 3 relates research questions to methods. For RQ1 we use a user survey (section 4). The survey results inform the remainder of the work. For RQ2 (section 5) and RQ3-5 (section 6) we apply methods from mining software repositories and software testing. We analyze GitHub issues (RQ2), correlate issue frequency to ONNX versions (RQ3), and evaluate ONNX on a range of model types (RQ4).
@@ 4. Theme 1: Interoperability Study
To understand the use of DL interoperability, we surveyed DL practitioners on DL development and deployment practices. We selected the survey methodology to gather insights across diverse practices and experiences (Easterbrook et al. 2008). Method is in section 4.1, results in section 4.2.
4.1. Methods
@@ 4.1.1. Instrument Design
We followed Kitchenham & Pfleeger’s guidelines to develop our survey instrument (Kitchenham and Pfleeger 2008). Table 2 illustrates the result. We asked general questions about DL interoperability, and specific questions about ONNX (interview data (Shankar et al. 2024) and GitHub suggest it is most popular). For higher response rate, the survey was short and mostly closed-ended. We iterated internally, piloting with 3 external participants to check instrument clarity.
@@ 4.1.2. Population and Sampling
With approval from our Institutional Review Board (IRB), survey respondents were recruited from Hugging Face users. The Hugging Face ecosystem is the primary location for deep neural network-based model development and re-use (Jiang et al. 2023d), so its users are a suitable population for questions about DL interoperability tools. To increase the likelihood of responses from experienced software engineers, we collected email addresses from users with PRO accounts (i.e., paid), and from accounts in organizations marked as company, community, and non-profit (excluding types such as education). Participants received $10 gift cards.
We targeted a confidence level of 90% with a 10% margin of error. With an estimated Hugging Face user population of 1.2 million (Forbes 2024; Gargi 2024), a sample size of 69 respondents was needed. We sent out surveys in batches until reaching our desired sample size. In total, we distributed our survey to 228 PRO users and 1,985 organization members. We received a total of 92 valid responses (4% response rate). All questions were optional to improve the response rate, so we do not have responses from all subjects on all questions.
@@
| Type | Break-down |
| ML Exp. | 1 yr. (8) ; 1-2 yr. (17) ; 3-5 yr. (32) ; 5 yr. (32) |
| SE Exp. | 1 yr. (4) ; 1-2 yr. (12) ; 3-5 yr. (20) ; 5 yr. (53) |
@@
| Deployment Env. | Web application (59) ; Cloud and data center (52) ; Desktop application (19) ; Mobile (14) ; IoT/embedded systems (14) ; Other (4) |
@@ 5. Theme 2: Failure Analysis (of ONNX)
The second theme of this work is a failure analysis of deep learning interoperability software, specifically, deep learning model converters. We apply the method of Failure Analysis (Petroski et al. 1994; Leveson 1995; Amusuo et al. 2022b; Anandayuvaraj and Davis 2022), which characterizes and reports the distributions of past failures to revise engineering methodologies and prioritize research targets.
@@
About ONNX: The ONNX (Open Neural Network eXchange) specification provides a common representation that DL frameworks and runtimes use to represent DL models (ONNX 2022). The ONNX specification has three main components: (1) A definition of a computational graph; (2) definitions of standard data types; and (3) definitions of built-in operators such as ArgMax (ONNX 2022). The ONNX specification changes regularly to keep up with DL framework evolution (Shridhar et al. 2020). These changes consist of adding new types and operators, and updating the behavior of existing operators. These changes are versioned within operator sets. As of 2023 there have been 18 operator sets, totaling 149 additions and 217 updates to the available operators.
5.1. Methods
@@ 5.1.1. Data Selection
| torch.onnx | PyTorch, Caffe2 | 77,429* | 20,938 | 1,286 (959)* |
@@
Using the PyTorch and TensorFlow converters covers two relevant differences within DL and interoperability. Within DL, these converters include both common representations of computational graphs: static (TensorFlow) and dynamic (PyTorch) (Hattori et al. 2020; Baydin et al. 2018). Within interoperability, these converters include both kinds of converter owners. Converters are naturally owned either by the upstream producer of the data or the downstream consumer of the data. In our case, torch.onnx gives an example of upstream ownership (it is owned by the PyTorch engineers), while tf2onnx gives an example of downstream ownership (it is owned by the ONNX engineers).
(2) Filters for relevance: For each repository, we collected closed GitHub issues related to ONNX conversion. For torch.onnx, the PyTorch repository marks ONNX converter issues with the label module: onnx. We collect all closed issues with this label, yielding the search filter: is:issue label:“module: onnx” is:closed. For tf2onnx, all issues are relevant. We remove issues labeled “pending on user response” and “question” because these issues may not be failures. This yields in 959 issues in torch.onnx (from 20,782 issues, not all ONNX-related) and 649 issues in tf2onnx (from 792 issues).
(3) Filters for data availability: Following prior work (Islam et al. 2019b; Franco et al. 2017; Garcia et al. 2020), we subsequently filtered for GitHub issues that contained enough information for failure analysis. We filter issues for sufficient information (e.g., the issue is resolved with a commit and pull request). This filtering is conducted upon the timeline events for each GitHub issue and the filtering criteria are given in Figure 4. This filter yields 327 issues in torch.onnx and 242 issues in tf2onnx.
@@ 5.1.2. Data Analysis
Symptoms & Causes: We adapted taxonomies from Shen et al. (Shen et al. 2021a), the closest related work. As that work concerns DL compilers rather than model converters, we trialed the taxonomies on a sample of 30 randomly chosen issues from each repository. Two researchers classified symptoms and causes. Computing inter-rater reliability via the Kappa coefficient (Cohen 1960), we found =0.90 (causes) and =0.95 (symptoms), i.e., “strong agreement” (McHugh 2012) for both taxonomies.
@@
Interrater Agreement: In our pilot study of applying Shen et al.’s taxonomies, we observed strong agreement between raters. Therefore, we relied on a single rating (one rater analyzing all issues) of the 200 sampled issues. To assess the risk of taxonomic drift (Wicks 2017), a second rater analyzed one tranche of randomly selected failures. In those 20 samples, the two raters had perfect agreement/=1.0. We conclude the single-rater analysis was sound.
@@ 5.2.2. Failure Symptoms
Our results show the similarity of the failure symptom distribution between DL model converters (our work) and DL compilers (Shen et al. (Shen et al. 2021a)). In the last two columns of Table 9, we see the Crash and Wrong Model symptoms make up the majority of symptoms across both DL converters and DL compiler front-ends. The Wrong Model and Crash symptoms appear characteristic of interoperability in this context, regardless of implementation.
| Symptom | TF | PT | Total | DL Comp. (Shen et al. 2021a) |

### v4

# Interoperability in Deep Learning: A User Survey and Failure Analysis of ONNX Model Converters

@@ Abstract.
This paper analyzes failures in DL model converters. We survey software engineers about DL interoperability tools, use cases, and pain points (N=92). Then, we characterize failures in model converters associated with the main interoperability tool, ONNX (N=200 issues in PyTorch and TensorFlow). Finally, we formulate and test two hypotheses about structural causes for the failures we studied. We find that the node conversion stage of a model converter accounts for 75% of the defects, and that 33% of reported failure are related to semantically incorrect models. The cause of semantically incorrect models is elusive, but models with behaviour inconsistencies share operator sequences. Our results motivate future research on making DL interoperability software simpler to maintain, extend, and validate. Research into behavioural tolerances and architectural coverage metrics could be fruitful.
@@ 1. Introduction
Deep Learning (DL) achieves state-of-the-art performance in many domains (Grigorescu et al., 2020; Kim et al., 2019). Software engineers engage in many activities for deep learning, including developing, re-using, fine-tuning, and deploying DL models (Amershi et al., 2019; Han et al., 2021; Jiang et al., 2023, 2024a). They use tools at each stage: DL frameworks for development (e.g., PyTorch (Paszke et al., 2019)); DL model registries for re-use (e.g., HuggingFace (Face, 2021)), and DL compilers for deployment platforms (e.g., TVM (Chen et al., 2018)). Preferably, these tools would be interoperable, so that DL models can move seamlessly from one to another. Model conversion errors disrupt engineering workflows or compromise the resulting models (Openja et al., 2022). High-quality model converters are crucial to the deep learning ecosystem.
Researchers have characterized failures in most of the DL ecosystem, but not in model converters. As depicted in Figure 1, previous works have considered DL development frameworks (Long and Chen, 2022a, b; Islam et al., 2019; Pham et al., 2020; Tan et al., 2022) and DL deployment compilers and runtimes (Shen et al., 2021; Guo et al., 2020). In contrast, prior research on DL model converters is limited to measuring conversions of 5 DL models (Openja et al., 2022). We lack systematic knowledge of failure symptoms, causes, and patterns in DL model converters.
In this work, we analyze failures of DL model converters. We first survey software engineers (N=92), focusing on their experiences with DL interoperability tools, their use cases, and failure modes encountered. Then we analyze failures in DL model converters to the ONNX IR (Open Neural Network eXchange’s Intermediate Representation), the most prominent interoperability target for DL models. We sample 200 closed GitHub issues (100 per converter) and determine failure symptoms, causes, and locations. Finally, we examine two possible root causes of model converter failures, testing hypotheses about specification updates and model types.
@@ 2.1. DL Model Conversion as Interoperability
In the context of DL, interoperability focuses on model reuse and is defined as the ability of DL software to exchange DL models (i.e., deep neural networks/DNNs) (Liu et al., 2020). Wegner describes two patterns for interoperability: creating pairwise mappings between systems; and introducing a common intermediary understood by all participants (Wegner, 1996). These patterns are illustrated in Figure 1 in the context of DL. The two patterns trade-off customizability for scalability.
Two kinds of DL systems interoperate on DL models: frameworks for DL model development, and runtimes for DL model deployment. In the downward path of Figure 1, pairwise mappings occur in DL compilers such as TVM (Chen et al., 2018), which map from framework representations into internal compiler representations for each supported framework. Along the rightward path, common intermediaries such as ONNX (ONN, 2019) and MMDnn (Liu et al., 2020) give standard representations for DL models. In this pattern, each framework and runtime has a one- or two-way adapter to the common intermediary.
DL Model Converters: Model converters fill a purpose similar to compiler front-ends (Li et al., 2020). They transform a model from a DL framework into a high-level IR representing the model’s computations and control flow. Graph-level optimizations are applied before further conversion to a low-level IR for hardware optimization and code generation. Table 1 summarizes a typical design.
@@
Figure 2 illustrates a model converted from PyTorch to ONNX. Conversion is challenging, as noted by AirBus (ONNX, 2023; Gauffriau et al., 2024) and others (Consortium, 2023), because it maps between graphs expressed with different operators and different semantics. Model conversion can produce models that are incompatible with runtimes or have different behaviours. For a compatibility issue, in PyTorch #78721 a converted ONNX model had a type mismatch (vbogach, 2022). For a behavioral issue, in PyTorch #74732 a converted ONNX model’s prediction changed (YsYusaito, 2022).
After conversion to an IR, models can be rendered back to DL frameworks or deployed to hardware. Framework-to-framework conversion can bypass issues in model reengineering (Jiang et al., 2024a). Deployment from an IR allows DL runtimes to optimize against the IR rather than the many DL frameworks.
@@ 2.2. Failure Studies of DL Components
Software engineering failures inform development and maintenance (Petroski et al., 1994; Anandayuvaraj and Davis, 2023; Amusuo et al., 2022). Previous failure analyses of DL interoperability software focused on the “Development” and “Runtime” components of Figure 1. Chen et al. studied 800 defects from 4 DL frameworks to obtain testing guidelines (Long and Chen, 2022a). Shen et al. studied DL compiler defects (Shen et al., 2021), and others have tested DL compilers (Liu et al., 2023; Xiao et al., 2022). Some of this work has incidentally examined interoperability failures (Liu et al., 2023; Morovati et al., 2024); for example, Shen et al.’s study of DL compilers included “front ends” offering interoperability through pairwise (DL framework specific) and common intermediary (ONNX) model loaders (Shen et al., 2021);
The two prior failure studies of DL model converters focused on testing and fault localization. Openja et al. converted 5 popular DL models (Openja et al., 2022). They considered only the current state of model converters, not past failures. Louloudakis et al. studied behavioral issues resulting from framework-to-framework conversion (Louloudakis et al., 2023a). They found failures in 10 out of 36 conversions. They created a fault localization and repair pipeline to localize and fix discrepancies (Louloudakis et al., 2023b).
@@ 3.2. Study Design
Figure 3 relates research questions to methods. For RQ1 we use a user survey (section 4). The survey results inform the remainder of the work. For RQ2 (section 5) and RQ3-4 (section 6) we apply methods from mining software repositories and software testing. We analyze GitHub issues (RQ2), correlate issue frequency to ONNX versions (RQ3), and evaluate ONNX on a range of model types (RQ4).
@@ 4. Theme 1: Interoperability Study
To understand the use of DL interoperability, we surveyed DL practitioners on DL development and deployment practices. We selected the survey methodology to gather insights across diverse practices and experiences (Easterbrook et al., 2008). Our method is given in section 4.1, results in section 4.2.
4.1. Methodology
@@ 4.1.1. Instrument Design
We followed Kitchenham & Pfleeger’s guidelines to develop our survey instrument (Kitchenham and Pfleeger, 2008). Table 2 illustrates the result. We asked general questions about DL interoperability, and specific questions about ONNX (interview data (Shankar et al., 2024) and GitHub suggest it is most popular). We removed some demographic questions after the pilot study (e.g., ML/SE expertise, organization types) to reduce survey time in order to increase response rates, as is recommended by survey guidelines such as (Ghazi et al., 2019). The pruning removed both demographic and technical questions, and reduced the survey time from 15-20 minutes (with which we had a 1% response rate from 130 initial invitations) to 5-8 minutes (yielding a 4% response rate from subsequent invitations). We made sure that we did not prune too far, so that the collected demographic data (reported in Table 3) is consistent with prior empirical software engineering works such as (Jiang et al., 2023; Michael et al., 2019; Ernst et al., 2015). We iterated internally, piloting with 3 external participants to check instrument clarity.
@@ 4.1.2. Population and Sampling
With approval from our Institutional Review Board (IRB), survey respondents were recruited from Hugging Face users. The Hugging Face ecosystem is the primary location for deep neural network-based model development and re-use (Jiang et al., 2023; Jones et al., 2024), so its users are a suitable population for questions about DL interoperability tools. To increase the likelihood of responses from experienced software engineers, we collected email addresses from users with PRO accounts (i.e., paid), and from accounts in organizations marked as company, community, and non-profit (excluding types such as education). Participants received $10 gift cards.
We targeted a confidence level of 90% with a 10% margin of error. With an estimated Hugging Face user population of 1.2 million (Forbes, 2024; Gargi, 2024), a sample size of 69 respondents was needed. We sent out surveys in batches until reaching our desired sample size. In total, we distributed our survey to 228 PRO users and 1,985 organization members. We received a total of 92 valid responses (4% response rate). All questions were optional to improve the response rate, so we do not have responses from all subjects on all questions.
@@
| Kind | Distribution of Responses |
| ML experience | 1 yr. (8) ; 1-2 yr. (17) ; 3-5 yr. (32) ; 5 yr. (32) |
| SE experience | 1 yr. (4) ; 1-2 yr. (12) ; 3-5 yr. (20) ; 5 yr. (53) |
@@
| Deployment environment | Web application (59) ; Cloud and data center (52) ; Desktop application (19) ; Mobile (14) ; IoT/embedded systems (14) ; Other (4) |
@@ 5. Theme 2: Failure Analysis (of ONNX)
The second theme of this work is a failure analysis of deep learning interoperability software, specifically, deep learning model converters. We apply the method of Failure Analysis (Petroski et al., 1994; Leveson, 1995; Amusuo et al., 2022; Anandayuvaraj and Davis, 2023), which characterizes and reports the distributions of past failures to revise engineering methodologies and prioritize research targets.
@@
About ONNX: The ONNX (Open Neural Network eXchange) specification provides a common representation that DL frameworks and runtimes use to represent DL models (ONNX, 2022). The ONNX specification has three main components: (1) A definition of a computational graph; (2) definitions of standard data types; and (3) definitions of built-in operators such as ArgMax (ONNX, 2022). The ONNX specification changes regularly to keep up with DL framework evolution (Shridhar et al., 2020). These changes consist of adding new types and operators, and updating the behavior of existing operators. These changes are versioned within operator sets. As of 2023 there have been 18 operator sets, totaling 149 additions and 217 updates to the available operators.
5.1. Methodology
@@ 5.1.1. Data Selection
| torch.onnx | PyTorch, Caffe2 | *77,429 | 20,938 | *1,286 (959) |
@@
Using the PyTorch and TensorFlow converters covers two relevant differences within DL and interoperability. Within DL, these converters include both common representations of computational graphs: static (TensorFlow) and dynamic (PyTorch) (Hattori et al., 2020; Baydin et al., 2018). Within interoperability, these converters include both kinds of converter owners. Converters are naturally owned either by the upstream producer of the data or the downstream consumer of the data. In our case, torch.onnx gives an example of upstream ownership (it is owned by the PyTorch engineers), while tf2onnx gives an example of downstream ownership (it is owned by the ONNX engineers).
(2) Filters for Relevance: We follow prior work (Garcia et al., 2020; Romano et al., 2022; Chaliasos et al., 2021), and study closed issues because these issues typically contain greater information about failure causes and symptoms. For each repository, we collected closed GitHub issues related to ONNX conversion. For torch.onnx, the PyTorch repository marks ONNX converter issues with the label module: onnx. We collect all closed issues with this label, yielding the search filter: is:issue label:“module: onnx” is:closed. For tf2onnx, all issues are relevant. We remove issues labeled “pending on user response” and “question” because these issues may not be failures. This yields in 959 issues in torch.onnx (from 20,782 issues, not all ONNX-related) and 649 issues in tf2onnx (from 792 issues).
(3) Filters for Data Availability: Following prior work (Islam et al., 2019; Garcia et al., 2020), we subsequently filtered for GitHub issues that contained enough information for failure analysis. We filter issues for sufficient information (e.g., the issue is resolved with a commit and pull request). This filtering is conducted upon the timeline events for each GitHub issue and the filtering criteria are given in Figure 4. This filter yields 327 issues in torch.onnx and 242 issues in tf2onnx.
@@ 5.1.2. Data Analysis
Symptoms & Causes: We adapted taxonomies from Shen et al. (Shen et al., 2021), the closest related work, as discussed in section 2.2. Conceptually, we chose to reuse deep learning compiler taxonomies due to the expected similarity between compiler front-ends and converters. This choice also follows the recent advice of Amusuo et al., who called for greater reuse of taxonomies in failure analysis research, where feasible (Amusuo et al., 2022). To assess whether the taxonomies are applicable to converters, we trialed the taxonomies on a sample of 30 randomly chosen issues from each repository. Two researchers classified symptoms and causes. Computing inter-rater reliability via the Kappa coefficient (Cohen, 1960), we found =0.90 (causes) and =0.95 (symptoms), i.e., “strong agreement” (McHugh, 2012) for both taxonomies.
@@
Interrater Agreement: In our pilot study of applying Shen et al.’s taxonomies, we observed strong agreement between raters. Therefore, we relied on a single rating (one rater analyzing all issues) of the 200 sampled issues. To assess the risk of taxonomic drift (Wicks, 2017), a second rater analyzed one tranche of randomly selected failures. In those 20 samples, the two raters had perfect agreement/=1.0. We conclude the single-rater analysis was sound.
@@ 5.2.2. Failure Symptoms
Our results show the similarity of the failure symptom distribution between DL model converters (our work) and DL compilers (Shen et al. (Shen et al., 2021)). In the last two columns of Table 9, we see the Crash and Wrong Model symptoms make up the majority of symptoms across both DL converters and DL compiler front-ends. The Wrong Model and Crash symptoms appear characteristic of interoperability in this context, regardless of implementation.
| Symptom | TF | PT | Total | DL Comp. (Shen et al., 2021) |